OpenStudy (anonymous):

evaluate each expression. LOGS , CLICK 2 SEE.

OpenStudy (anonymous):

\[Log _{125}5\]

OpenStudy (anonymous):

alright 125^x=5

OpenStudy (anonymous):

it is another way to write a log

OpenStudy (ksaimouli):

@cherio12 that is not =x or something that = or equation is wrong

OpenStudy (ksaimouli):

1/3

OpenStudy (anonymous):

what do you mean ksaimouli?

OpenStudy (anonymous):

\[125^{x}=5\] so its not this ?

OpenStudy (anonymous):

no it is...i mean if you do 125^(1/3) you get 5

OpenStudy (anonymous):

i just like rewritting it that way to help me think about it

OpenStudy (anonymous):

so what steps would you take to get that answer ? how would you completely solve it. just that one step ? rewriting the equation ? @cherio12

OpenStudy (anonymous):

yes, that it all the work i would show

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

after writing 125^x= 5 how do you find that x = 1/3?

OpenStudy (anonymous):

the way i think about it is 125 to the what would give me 5.

OpenStudy (anonymous):

it is helpful to know the first couple power of numbers. 5's are 5, 25, 125, 625..

OpenStudy (anonymous):

it just comes with time